Mr. Ahmed is a sales manager at the mobile company. His salary includes commission, and basic salary. His basic salary is Rs. 25,000 per month and he is given Rs. 325 for the sale of each mobile. If his total salary for last month was Rs. 97,475. How many mobiles did he sell last month?

1
Expert's answer
2022-03-22T18:58:30-0400

The salary is "25000+325n" , where n is the number of phones sold.

"25000+325n=97475\Rightarrow n=\frac{97475-25000}{325}=223"